OpenStudy (anonymous):

Write the function that models each relationship. Find z when x=4 and y=8. z varies jointly with x and y. When x=2 and y=2, z=7.

OpenStudy (anonymous):

@xoxoLoveBre

OpenStudy (anonymous):

well it could be z=2(x+y)-1 check z=2(2+2)-1=7 z=2(4+8)-1=23